OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle you must show all work and calculation to recieve credit x^2 -4x+y^2+8y=-4

OpenStudy (anonymous):

I need help urgent

OpenStudy (zehanz):

Are you familiar with the concept of completing the square?

OpenStudy (anonymous):

Kind of

OpenStudy (zehanz):

Do you know the general form of the equation of a circle with radius r and center (a, b)?

OpenStudy (anonymous):

Yes

OpenStudy (zehanz):

So you know you have to rewrite the given equation to the form: (x-a)²+(y-b)²=r². First, try to complete the square in x² - 4x.

OpenStudy (zehanz):

Sorry, that looks like incomprehensible stuff!

OpenStudy (anonymous):

Sorry i am on a ipad and cannot see that its just a bunch of squared

OpenStudy (zehanz):

So you know you have to rewrite the given equation to the form: (x-a)^2+(y-b)^2=r². First, try to complete the square in x^2 - 4x.

OpenStudy (zehanz):

It should read r^2 after the =

OpenStudy (anonymous):

I do not understand how would i complete theat part of the equation

OpenStudy (zehanz):

you could write x^2-4x = (x-2)^2 - 4, because, if you expand this, you''l get: x^2-4x+4-4, which is what you had at the beginning. This is called "completing the square". So, in your equation, you can replace the x^2-4x with : (x-2)^2 - 4.

OpenStudy (zehanz):

It has to do with this well-known rule: (a - b)^2 = a^2 - 2ab + b^2

OpenStudy (anonymous):

Ok how do we continue

OpenStudy (zehanz):

Do the same with y^2 + 8y

OpenStudy (anonymous):

(Y-2)^4???

OpenStudy (zehanz):

No, just make a complete square! If you take (y +4)^2 and expand it, you have: y^2 + 8y + 16. This is almost what is in your equation: there it reads: y^2 + 8y. So if you would replace y^2 + 8y with (y + 4)^2, it would be 16 to big. Well, then replace y^2 +8y with (y + 4)^2 - 16, then everybody is happy. We have completed the square.

OpenStudy (zehanz):

Do you understand what I did there?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you

OpenStudy (zehanz):

OK, fine! Now let us see what we have: x^2 -4x+y^2+8y=-4 was the original equation. we replaced x^2 - 4x with (x - 2)^2 -4. we replaced y^2 + 8y with (y +4)^2 - 16. This leaves us with: (x-2)^2 - 4 + (y + 4)^2 - 16 = -4. Now for the last step: convert to the form: (x - a)^2 + (y - b)^2 = r^2. Can you do this?

OpenStudy (anonymous):

Not sure but (x+6)^2+(y-12)^2=r^2

OpenStudy (zehanz):

You are confusing the squares and the constants. The (x-2)^2 and (y + 4)^2 parts are the squares. You cannot just add some constants like -4 and -16 to the stuff inside the brackets, because of the ^2. But there is no need to do that. See the -4 on both sides of the equation? Add 4 on both sides and poof! the're gone, leaving us with (x -2)^2 + (y + 4)^2 - 16 = 0. Now the general form is (x - a)^2 + (y - b)^2 = r^2. If we add 16 on both sides we have: (x -2)^2 + (y + 4)^2 = 16, or: (x -2)^2 + (y + 4)^2 = 4^2. Now we can read off the midpoint: it is (2, -4), because a = 2 and b = -4. Also r = 4. We're done! Advice: look up the theory on this, because I have the feeling you do not understand this procedure completely...

OpenStudy (anonymous):

Okay i will look up that info thank you for all your help iwould probably still be stuck on this right now. Thanks again

OpenStudy (zehanz):

yw!